log3 (x+2) + log3 (x-2) = 1
Answer by stanbon(29603)

log3 (x+2) + log3 (x-2) = 1
log[(x+2)(x-2)] = 1
(x+2)(x-2) = 3^1
x^2-4 = 3
x^2-7 = 0
(x-sqrt7)(x+sqrt7) = 0
x = sqrt7 or x=-sqrt7
=======================
Cheers,
Stan H.